A nozzle is a device for increasing the velocity of a steadily flowing stream. At the inlet to a certain nozzle, the enthalpy of the fluid passing is 3025 kJ/kg and the velocity is 60 m/s. At the discharge end, the enthalpy is 2790 kJ/kg. The nozzle is horizontal and there is negligible heat loss from it.
a)Find the velocity of the fluid at the exit.

(b the rate of flow of fluid when the inlet area is 0.1 m^2 and the specific volume at inlet is 0.19 m^3/kg,

(c) the exit area of the nozzle when the specific volume at the nozzle exit is 0.5 m^3/kg
1
Expert's answer
2020-03-23T12:57:53-0400

a) From the conservation of energy:


"h _\n1\n\u200b\t\n \u200b+0.5v _\n1\n^2\n\u200b\t\n \u200b=h _\n2\n\u200b\t\n \u200b+0.5v_\n2^\n2\n\u200b\t\n \u200b"

"3025\u22c510^\n3\n +0.5(60) ^\n2\n =2790\u22c510 ^\n3\n +0.5v _\n2^\n2\n\u200b\t\n \u200b"

"v_2=688.2\frac{m}{s}"

b)


"\frac{dm}{dt}=\frac{\frac{dV}{dt}}{\frac{dV}{dm}}"

"\frac{dV}{dt}=Av_1"

"\frac{dm}{dt}=\frac{60\cdot 0.1}{0.19}=31.6\frac{kg}{s}"

c)


"\frac{dm}{dt}=\frac{v_2A'}{\frac{dV'}{dm}}"

"31.6=\frac{688.2A'}{0.5}"

"A'=0.0229\ m^2"